OpenStudy (anonymous):

What rate of interest, to the nearest tenth of a percent, compounded annually is needed for an investment of $240 to grow to $350 in 5 years?

OpenStudy (anonymous):

solve \[350=240(1+r)^5\] for r

OpenStudy (anonymous):

divide by 240, take the fifth root (with a calculator), then subtract 1

OpenStudy (anonymous):

\[\frac{350}{240}=(1+r)^5\] \[1+r=\sqrt[5]{\frac{350}{240}}\] i get 1.0784

OpenStudy (anonymous):

so \[r=.0784=7.84\%\] rounded of course

OpenStudy (anonymous):

thanks so much! I kept getting stuck in the beginning!